OpenStudy (anonymous):

Need help understanding the concept Let G1 = , G2 = 0, x(multiplication)> f:G1--> G2 defined by f(x) = e^x Are G1, G2 isomorphic? the answer is yes and the logic is \(e^{x+y}=e^x *e^y\). This preserves groups structure, preserves multiplication table. Please, explain me what it mean.

OpenStudy (anonymous):

\(G_1=\langle\mathbb{R},+\rangle\) and \(G_2=\langle \mathbb{R}_{+},\times\rangle\) ? where \(\mathbb{R}_+\) is positive reals only?

OpenStudy (anonymous):

Yes, Sir.

OpenStudy (anonymous):

Okay, in this case, an isomorphism is a bijective function \(f:G_1\to G_2\) such that \(f(u+v)=f(u)\times f(g)\) for all \(u,v\in G_1\), which is the case for the exponential function.

OpenStudy (anonymous):

\[f(x+y)=e^{x+y}\] \[f(x)\times f(y)=e^x\times e^y\]

OpenStudy (anonymous):

Yes, I got it so far, but "it preserves groups structure" mean??

OpenStudy (anonymous):

* the \(f(g)\) should say \(f(v)\), sorry.

OpenStudy (anonymous):

Well, what is a group? It's a set with an operation that is closed under the operation, associative, has an identity element, and is invertible. \(f\) is closed under the operation, since \(e^x\) and \(e^y\) are real, and so is the product \(e^{x+y}\). \(f\) is associative, since \(e^{(x+y)+z}=e^{x+(y+z)}\) and \((e^xe^y)e^z=e^x(e^ye^z)\). The identity element is 0, since \(e^{0+x}=e^0e^x=e^x\). I'm not sure about invertibility, but I think it has to do with the logarithm function.

OpenStudy (anonymous):

You mean after working on the function f, the range of f = set of elements in G2, right?

OpenStudy (anonymous):

Actually, that identity element might not be right... which means the inverse element has less to do with logs. Yes, the range is the set associate with \(G_2\).

OpenStudy (anonymous):

I think I got it. Thanks a lot, Professor. :)

OpenStudy (anonymous):

I watch this video, this problem is at 21:23 https://www.youtube.com/watch?v=mwcNETa0KFI

OpenStudy (anonymous):

Maybe the identity is just 1, which makes the inverse element \(\dfrac{1}{x}\) for \(x\in G_1\), \(x\not=0\). You don't have to worry about that for \(x\in G_2\) though.

OpenStudy (anonymous):

And you're welcome, but I'm no professor :P

OpenStudy (anonymous):

Ok, I remember that once you said "I..... , I am bad professor" . That's why I think you are. :)

OpenStudy (anonymous):

Hmm, I don't remember saying anything like that... Oh well, happy to help :)

OpenStudy (anonymous):

:)